square root 2x-1 +8=x
please help!!!

Answer by edjones(7311)

sqrt(2x-1)+8=x I assume that this is the problem. Use parentheses!
sqrt(2x-1)=x-8
2x-1=x^2-16x+64 square each side.
x^2-18x+65=0
(x-13)(x-5)=0 Factor
x=13, x=5
.
Ed